OpenStudy (anonymous):

simplify (square root 3) (5 - square root 6)

OpenStudy (anonymous):

5sqrt(3) -sqrt(18) = 5sqrt(3) - 3sqrt(2)

OpenStudy (anonymous):

\[5\sqrt{3}-\sqrt{3}\sqrt{6}\]\[5\sqrt{3}-3\sqrt{2}\]

OpenStudy (anonymous):

Clarification:\[\sqrt{3}\sqrt{6}=\sqrt{3\times6}\]\[\sqrt{3\times6}=\sqrt{9\times2}\]\[\sqrt{9\times2}=\sqrt{9}\sqrt{2}\]\[\sqrt{9}\sqrt{2}=3\sqrt2\]